3.A quantity with an initial value of 980 grows exponentially at a rate of 9.5% every 7 years. What is the value of the quantity after 93 years, to the nearest hundredth?


1
Expert's answer
2021-01-27T09:53:32-0500

We have the following quantities

"a = 980, \space\space r = 0.095, \space\space \space\space \space\space"

"\gamma - ? \space\space t = 93years = \frac{93}{7} = 13.28",

"\gamma = a*(1+r)^t = 980 * (1+0.095)^{13.28},\\\space\space γ=a∗(1+r)^ 
t = 3270"